Solve the following differential equation.
(3x +2y +y2)dx+(x+4xy+5y2)dy = 0.
Help please
1
Expert's answer
2021-08-27T10:01:52-0400

Solution

Check the exactness of the equation:

"M=3x+2y+y^2"

"N=x+4xy+5y^2"

"\frac{dM}{dy}=2+2y"

"\frac{dN}{dx}=1+4y"

Clearly,

"\frac{dM}{dy}\neq\frac{dN}{dx}"

The equation is not exact.

Take "4x+4y^2" as the integrating factor;

"M'=(3x+2y+y^2)(4x+4y^2)"

"M'=12x^2+8xy+16xy^2+8y^3+4y^4"

Hence

"\frac{dM'}{dy}=8x+32xy+24y^2+16y^3"

"N'=(x+4xy+5y^2)(4x+4y^2)"

"N'=4x^2+16x^2y+24xy^2+16xy^3+20y^4"

"\frac{dN'}{dx}=8x+32xy+24y^2+16y^3"

in which

"\frac{dM'}{dy}=\frac{dN'}{dx}"

The equation is now exact.

The general solution is

"\int M'dx+\int(" Terms in N independent of x")dy=C"

"\int(12x^2+8xy+16xy^2+8y^3+4y^4)dx+\int20y^4dy=C"

"=4x^3+4x^2y+8x^2y^2+8xy^3+4xy^4+4y^5=C"